OpenStudy (anonymous):

the integral of (sqrt of x - 1/sqrt of x) dx

OpenStudy (anonymous):

\[\int\limits_{}^{} (\sqrt{x}-1/\sqrt{x})\]

OpenStudy (anonymous):

Rewrite it as: \[\int\limits x^{\frac{1}{2}}-x^{\frac{-1}{2}}dx\] Then use the power rule as before. :)

OpenStudy (anonymous):

oh. now that made the equation easier to solve. lol. thanks! :)

OpenStudy (anonymous):

No problem :P